If 𝑓(𝑥)=1𝑥, show that 𝑓(𝑎)−𝑓(𝑏)=𝑓(𝑎𝑏/𝑏−𝑎).


1
Expert's answer
2021-10-20T01:15:00-0400

Solution:

Given "f(x)=\dfrac1x"

"LHS=f(a)-f(b)
\\=\dfrac1a-\dfrac1b
\\=\dfrac{b-a}{ab}"

"RHS=f(\dfrac{ab}{b-a})
\\=\dfrac{1}{\dfrac{ab}{b-a}}
\\=\dfrac{b-a}{ab}"

So, LHS=RHS

Hence, proved.